Given 𝑓(π‘₯) = π‘Žπ‘₯3 + 2π‘₯2 + 𝑏π‘₯ βˆ’ 20 where π‘₯ βˆ’ 2 and π‘₯ βˆ’ 5 are two of its factors determine the real values of π‘Ž and 𝑏.Β 


1
Expert's answer
2021-11-07T17:27:02-0500

f(x) = ax3+2x2+bx-20


Let's find value of f(x) at x = 2, x = 5

f(2) = a*(2)3+2*(2)2+b*2-20

f(2) = 4a + b - 6


f(5) = a*(5)3+2*(5)2+b*5-20

f(5) = 25a + b + 6


Let's add f(2) + f(5) and solve for b

4a + b - 6 + 25a + b + 6 = 29a + 2b

29a +2b = 0

b = -29a / 2


Let's subtract f(2) - f(5) to find a

4a + b - 6 - (25a + b + 6) = -21a - 12

-21a - 12 = 0

a = 12 / (-21)

a = -(4/7)


find b

b = (-29 * (-4/7)) / 2

b = 58/7


Answer: a = - 4/7, b = 58/7